OpenStudy (anonymous):

How many total atoms are in 0.240 g of P2O5?

OpenStudy (abb0t):

0.240 g \(\times\) molar mass of P\(_2\)O\(_5\) \(\times\) \(\sf \color{red}{6.022 \times 10^{23} \frac{atoms}{mol}}\)

OpenStudy (anonymous):

how do you find the molar mass?

OpenStudy (abb0t):

Look at your periodic table. It is the total mass of each atoms times the number of atoms. So you have 2\(\times\)(mass of \(\sf \color{red}{P}\)) + 5 \(\times\) (mass of \(\sf \color{purple}{O}\)) = molar mass \(\sf \color{orange}{\frac{grams}{mol}}\)

OpenStudy (abb0t):

The mass is the number under the elemnt.t